OpenStudy (anonymous):

How do I do this type of questions? Calc 1

OpenStudy (anonymous):

It says, find f '(x), if\[\large f(x)=\int\limits_{4x}^{1}~\cot^2(t)~dx\] Can someone guide me through the problem please?

OpenStudy (anonymous):

@wio , can you help please? sorry for bothering you....

OpenStudy (anonymous):

I know how to integrate cot^2t, but what do I have to do, can you guide me please?

OpenStudy (anonymous):

Well, suppose that: \[ g'(t) = \cos^2(t) \]

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

That would mean that: \[ f(x) = \int_{4x}^1g'(t)~dt=g(1)-g(4x) \]

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

Therefore: \[ f'(x) = \left(g(1)-g(4x)\right)' \]

OpenStudy (anonymous):

Can you do it now?

OpenStudy (anonymous):

Let's see, so first I integrate.... wait, I don't have to integrate, I can tell it is -cot^2(4x), right?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

because the g(1) is constant, so the derivative will make it a zero.

OpenStudy (anonymous):

We don't need to find \(g(t)\).

OpenStudy (anonymous):

I have to integrate, yes?

OpenStudy (anonymous):

No, you don't need to integrate. You need to differentiate.

OpenStudy (anonymous):

find the derivative of cot^2(t) ?

OpenStudy (anonymous):

Do you need help differentiating here? \[ f'(x) = \left(g(1)-g(4x)\right)' \]

OpenStudy (anonymous):

I don't understand your set up so well.

OpenStudy (anonymous):

Does this make sense? \[ f'(x) = \frac{d}{dx}f(x) = \frac{d}{dx}\left(g(1)-g(4x)\right) \]

OpenStudy (anonymous):

\[\frac{ d }{ dx }\left(\begin{matrix} \cot^2(1)-\cot^2(4x)\end{matrix}\right)\]Like this?

OpenStudy (anonymous):

Okay, we don't know what \(g(t)\) is. We can simplify without knowing what it is.

OpenStudy (anonymous):

First step:\[ \frac{d}{dx}\left(g(1)-g(4x)\right) = \frac{d}{dx} g(1) - \frac{d}{dx}g(4x) \]

OpenStudy (anonymous):

Next step:\[ = - \frac{d}{dx}g(4x) \]

OpenStudy (anonymous):

So my set up, \[\frac{ d }{ dx }\left(\begin{matrix} \cot^2(1)-\cot^2(4x)\end{matrix}\right)\] was correct?

OpenStudy (anonymous):

It was wrong... that's what I've been saying.

OpenStudy (anonymous):

Sorry

OpenStudy (anonymous):

I am stupid.

OpenStudy (anonymous):

Do you know how you would simplify this:\[ - \frac{d}{dx}g(4x) \]

OpenStudy (anonymous):

IF I knew the g(x), what is it?

OpenStudy (anonymous):

You don't really need to know. Let's just say \(u=4x\) to make things simpler. Could you do this? \[ \frac{d}{dx}g(u) \]

OpenStudy (anonymous):

g'(u) ?

OpenStudy (anonymous):

No, you need to remember the chain rule.

OpenStudy (anonymous):

\[\frac{ dy }{ dx }=\frac{ dy }{ du }~\frac{ du }{ dx }\]this?

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

In our case: \[ \frac{d}{dx}g(u)=\frac{dg}{du} \frac{du}{dx} =g'(u) (4x)' = g'(4x)(4x)' \]

OpenStudy (anonymous):

Now, we said that: \[ g'(t) = \cot^2(t) \]

OpenStudy (anonymous):

That means that: \[ g'(4x)=\ldots \]

OpenStudy (anonymous):

cot^2(4x).... whoo

OpenStudy (anonymous):

the chain rule for the 4x....

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

\[ f'(x) =- 4\cot^2(4x) \]

OpenStudy (anonymous):

The \((4x)' = 4\).

OpenStudy (anonymous):

Yes, it is -, because it was g(1)-g(4x) Not like I am correcting you or anything like that. I think I kind of understand it, but still really poorly.

OpenStudy (anonymous):

I'll read some tutorials online and my notes, I guess I can better understand it then. Tnx for the time and effort.

OpenStudy (anonymous):

The only problem here is that you're not completely comfortable with your calculus as an algebraic operation.

OpenStudy (anonymous):

The implicit differentiation part.

OpenStudy (anonymous):

Perhaps, that is it. Like I know just how to "calculate" but for the rest my head is not working-:(

OpenStudy (anonymous):

It is just harder for me to understand the query of the querstion.... well, tnx again:)

OpenStudy (anonymous):

I'll do the problem once more quickly so you can review.

OpenStudy (anonymous):

that's very nice;)

OpenStudy (anonymous):

Here is the entire thing. You can look at each step and let me know if there is a step that you just don't get.

OpenStudy (anonymous):

Okay... will see.

OpenStudy (anonymous):

\[ \begin{split} f'(x) &= \frac{d}{dx}f(x)&\quad&\\ &= \frac{d}{dx}\int_{4x}^1\cot^2(t)~dt&&\text{let }g'(t) = \cot^2(t)\\ &= \frac{d}{dx}\int_{4x}^1g'(t)~dt\\ &= \frac{d}{dx}(g(1)- g(4x))\\ &= \frac{d}{dx}g(1)- \frac{d}{dx}g(4x)\\ &= - \frac{d}{dx}g(4x)&&\text{let }u=4x\\ &= - \frac{d}{dx}g(u)\\ &= - g'(u)u'(x)&&\text{let } \cot^2(t)=g'(t)\\ &= - \cot^2(u)\frac{d}{dx}u&&\text{let } 4x=u\\ &= - \cot^2(4x)\frac{d}{dx}4x\\ &= - 4\cot^2(4x) \end{split} \]

OpenStudy (anonymous):

ohh, so the "let y = 4x" shouldn't be there?

OpenStudy (anonymous):

ohh, you fixed to u=4x... I se

OpenStudy (anonymous):

The y is next to u key.

OpenStudy (anonymous):

Yes yes:)

OpenStudy (anonymous):

u'(x) is the chain rule. Okay I think I am getting it.

OpenStudy (anonymous):

and you mean let g'(u) = cos^2(u).... right?

OpenStudy (anonymous):

I mean 4th to the last line...

OpenStudy (anonymous):

Yes, it is sort of the same thing.

OpenStudy (anonymous):

Yes, ty, I get it. I don't promise that I'll be able to do another one like this on my own, but I appreciate your help!

OpenStudy (anonymous):

I'll take a note of it on my papaer.

OpenStudy (anonymous):

There is a general formula for these things.

OpenStudy (anonymous):

So this is all I should have done?

OpenStudy (anonymous):

\[ \frac{d}{dx}\int_{a(x)}^{b(x)} f'(t)~dt = f'(b(x))b'(x)-f'(a(x))a'(x) \]

OpenStudy (anonymous):

yeah... I msee what you changed.

OpenStudy (anonymous):

the derivative, chain rule for each, right? for f(b) and f(a).

OpenStudy (anonymous):

Well, the thing is, it isn't that complicated to begin with, it's just when I do it with such detail, then it becomes very long.

OpenStudy (anonymous):

Yeah... I mean I am just very bad at problems that aren't like "differentiate f(x)" you know what I mean( ?) I won't necessarily know the "steps" but I am a pretty good calculator.

OpenStudy (anonymous):

You could have said: \[ \frac{d}{dx}\int_{4x}^1\cot^2(t)~dt = \cot^2(1)(1)' - \cot^2(4x)(4x)' = -4\cot^2(4x) \]But you might not see where that comes from.

OpenStudy (anonymous):

I see why is that (sort of).

OpenStudy (anonymous):

yes, I think I got it....tnx